# Revisiting 360 Depth Estimation with PanoGabor: A New Fusion Perspective

### v4

@@ Abstract
Index Terms:
Monocular, 360°, Depth Estimation, Distortion, Gabor Transform.
@@ II-A Monocular 360 Depth Estimation
Panoramic distortion leads to notable performance degradation in perspective image-based depth estimation approaches. To this end, Zioulis $et~{}al.$ [15] highlighted the need to learn within the 360 domain, introducing the 3D60 dataset. Moreover, to mitigate distortion, researchers have adjusted the sampling positions of convolutional kernels for distortion awareness [16, 17, 18]. Some indoor monocular panoramic depth estimation methods [19, 20, 21] have developed various distortion-aware techniques to manage panoramic distortions, but most of these techniques are computationally inefficient. Recently, advanced Transformer-based architectures [22] have broadened the receptive field, enabling models to extract features from 360 images. However, the traditional self-attention mechanism struggles to handle the complex distortion distribution in 360 images [4]. Enhanced distortion-aware transformer architectures [4, 5, 6, 23, 24, 25], while effective, often come with high computational and memory costs. To simplify models, some approaches [26, 27, 28, 29, 30] advocate for estimating indoor depth from vertically compressed one-dimensional sequences, which leverages gravity alignment properties. However, this compression often fails to capture essential structural geometric information. To achieve a balanced trade-off between performance and efficiency, the mainstream approaches[31, 32, 7] still prefer to develop a more efficient fusion-based framework. However, these fusion methods neglect the reintroduced distortion during the fusion process. Besides, although many of these approaches claim state-of-the-art performance, comparisons are often complicated by variations in datasets, preprocessing methods, and evaluation strategies, making fair comparisons challenging. To this end, Elite360D [7] provides a solid baseline. Therefore, in this paper, we follow the Elite360D method [7] to conduct the experiments.
@@ II-B Gabor Filters in Image Processing
The Gabor filter is widely used in texture analysis, object detection, and face recognition due to its robustness to varying angles and scales. Li $et~{}al.$ [33] suggested enhancing face recognition by learning the Gabor wavelet’s covariance matrix. In fabric defect detection, Chen $et~{}al.$ [34] introduced the Genetic Algorithm Gabor Faster R-CNN (Faster GG R-CNN), embedding Gabor kernels to reduce texture interference. Andrey $et~{}al.$ [35] proposed GaborNet, which uses Gabor filters to improve convergence and reduce the complexity of training in image recognition. To extend the receptive fields and enhance the geometric depth cues, we introduce Gabor filters to analyze texture in the frequency domain. Considering the special characteristics of 360 images, we propose a distortion-aware Gabor-based approach to address the challenging distortion issues.
@@ III Methodology
In this section, we introduce PGFuse, its pipeline is illustrated in Fig. 2. It takes an ERP image as input and produces the corresponding predicted depth map. Our model follows a typical encoder-decoder scheme and consists of three main components: a feature extractor, a unidirectional fusion module, and an up-sampling decoder. To reduce the computational burden, we employ a shared weight backbone to extract two sets of distortion-free cubemap features: one from the original perspective and the other horizontally rotated by 45°. The two groups of cube features are employed to relieve the pixel-missing and discontinuous boundary problems inherent in the cubemap projection[29]. These features are spatially aligned, unified into the ERP format, and then inputted into the CS-UFM module.
In the fusion module, we design PanoGabor filters to effectively handle the reintroduced distortion and enhance depth cues by analyzing textures in the frequency domain. Specifically, we introduce a latitude-aware distortion representation method that approximates the cosine value of latitude. We then design a Distortion-Aware PanoGabor Filters Algorithm 1 that uses this approach to generate PanoGabor filters. These filters are further enclosed as a convolution layer and embedded in the CS-UFM module. Subsequently, the features processed by the CS-UFM module are fed into a multi-scale distortion-aware module (MS-DAM) [25] to integrate the shallow geometric features and the deep semantic features. These integrated features are then fed into the decoder, following a hierarchical architecture with multi-scale stages. Each stage consists of an upsampling layer and a convolution layer, followed by an ELU activation function. Finally, a 3×3 convolution layer is used to produce the final depth map. In addition, we project the ERP depth map onto a spherical surface and employ the Sobel operator to compute depth gradients in the nearest tangent patch, providing additional constraints and stabilizing the orientation sensitivity of the PanoGabor filters. The following sections provide detailed descriptions of the proposed modules.
@@ III-A Preliminary
The distortion factor $K(\phi)$ ranges from $1$ to +$\infty$+ ∞. At the equator, the projection corresponds to an identity mapping, implying that the distortion is zero at this latitude. To more accurately describe the variation in distortion across latitudes, we introduce a bias term, leading to the following expression for the distortion factor:
@@
Gabor Transform. When applied to image processing, the Gabor transform can provide useful spatial frequencies and orientations analysis of the images. By responding to specific spatial frequencies and orientations, the Gabor filters can effectively identify textures and detect edges, which are critical for depth estimation. The standard Gabor transform formula can be denoted as:
@@
6:  $\theta\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow$ concatenate
7:  Calculate Gabor parameters [^1] The parameters are set up according to [11], $\epsilon$ is generated randomly.:$\psi\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow\pi\epsilon$; $f\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow(\frac{\pi}{2})\cdot(\sqrt{2})^{\epsilon}$;$\sigma\in\mathbb{R}^{C_{in}\times C_{out}}\leftarrow\frac{\pi}{f+0.1}$
8:  $f\leftarrow f\cdot(1+K(\phi_{C}))$;
9:  $(y,x)\leftarrow$ meshgrid((1,0,1),(1,0,1))101101((-1,0,1),(-1,0,1))( ( - 1 , 0 , 1 ) , ( - 1 , 0 , 1 ) ); $x,y\in\mathbb{R}^{3\times 3}$
10:  Rotate $x$ and $y$:$x_{rot}\leftarrow x\cdot\cos(\theta)+y\cdot\sin(\theta)$;$y_{rot}\leftarrow-x\cdot\sin(\theta)+y\cdot\cos(\theta)$
11:  Calculate Gabor filter $g$:$g\leftarrow\exp\left(-0.5\cdot\left(\frac{x_{rot}^{2}+y_{rot}^{2}}{\sigma^{2}}\right)\right)\cdot\cos(f\cdot x_{rot}+\psi)$;$g\leftarrow\frac{g}{2\cdot\pi\cdot\sigma^{2}}$
@@ III-B Distortion-aware PanoGabor Filters (PGF)
where the range of $d_{c}$ is $[0,+\infty)$. The scale factor mirrors the increasing distortion in equirectangular projections towards the poles. Subsequently, as detailed in Algorithm 1, we employ the defined distortion coefficient to control the generation of distortion-aware PanoGabor filters, thereby creating a bank of PanoGabor filters.
@@
Figure 5: Illustration of the spherical tangent sampling. a represent that the pixels are projected from the pixel coordinates to the spherical coordinates to sampling tangent points. b indicates that the calculated sampling points are projected back to the pixel coordinates.
@@ III-C Spherical Gradient Constraint
Create sampling mesh grid. Given a depth map $D\in\mathbb{R}^{H\times W}$, we first define the sampling grid. For each pixel in the depth map, we sample a 3×3 tangent patch. Let $X=[-1,0,1]$ and $Y=[-1,0,1]$ represent the sampling kernels for the $x$- and $y$-directions, respectively. We define the 2D grid mesh $M$ (Fig. 5left) as:
@@ III-D Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM)
The fusion module is designed to integrate two sets of ERP features, addressing reintroduced distortion, inconsistencies in cubemap boundaries, and pixel loss issues. This module (as illustrated in Fig. 6) first concatenates the two groups of input ERP features and uses a residual block to achieve an initial refinement. In particular, the block comprises a 1x1 convolution that compresses the channels to match the features’ height, facilitating the application of the PanoGabor layer. Following the 1×1 convolution layer, a 3x3 PanoGabor layer is used to process the features, aiming for distortion awareness and depth cues enhancement.
@@ IV Experiment
We conduct all experiments on a single NVIDIA GeForce RTX 3090 GPU, implementing our approach using the PyTorch framework. We use the Adam optimizer [39] with an initialized learning rate of $1\times 10^{-4}$ to train our model. We keep the default settings for other optimizer hyper-parameters. During the training period, we use a batch size of 8 and employ left-right flipping, panoramic horizontal rotation, and luminance change for data augmentation, as previously proposed in works [2, 4]. The backbones we employed are all pre-trained on ImageNet-1K.
Dataset. Two real-world datasets for experimental validation: Stanford 2D-3D [12], and Matterport3D [40]. The depth maps in Stanford 2D-3D are constructed from reconstructed 3D models, and some depths are inaccurate. Besides, the bottom and top regions of the RGB images are missing. We follow the official splits and get 1,040 RGB-D pairs for training and 343 for testing. The Matterport3D dataset provides raw depth whose top and bottom depths are missed. For the RGB panoramas in Matterport3D, we follow the Unifuse method [41] to fill the missing regions. Finally, we get 7829 pairs for training and 2014 pairs for testing.
One virtue dataset: Structrued3D [14]. The virtue dataset was first proposed to explore the layout estimation task. Recently, some researchers have preferred to evaluate the depth estimation models on the dataset for its high-quality annotations. We follow the official splits that we employ the frame scene_0000 to frame scene_02999 for training and scene_03250 to scene_03499 for testing. We especially follow the official guidelines and remove invalid cases.
@@
Figure 7: Qualitative comparison results. We exhibit the predicted depth maps of the Unifuse [2] and ours. All depth maps adopt a unified visualization method.
Figure 8: Visualization of the point cloud generated with the ground truth (GT) and the predicted depth map (Ours and UniFuse). The areas marked in red rectangles indicate that our approach gains superior performance (dense and high-quality point cloud).
@@ IV-B Comparison Analysis
Quantitative Results. We exhibit the quantitative comparison results in Table LABEL:tab:quanti, which demonstrate that our model outperforms all other state-of-the-art solutions on most metrics. Our PGFuse method significantly outperforms compared methods. Specifically, on the real-world Matterport3D and Stanford2D3D datasets, our approach demonstrates notable improvements in the important RMSE and $\delta_{1}$ metrics. Additionally, it achieves comparable results on the virtual Structured3D dataset, showcasing its robustness and effectiveness across various environments. These results highlight the efficacy of PGFuse in monocular 360-depth estimation across diverse and challenging datasets. Although our performance on the loosest accuracy metric $\delta_{3}$ is slightly lower than the Elite360D [7] and the Unifuse [2], it performs better than the two methods on the other two tighter accuracy metrics ($\delta_{1}$ and $\delta_{2}$). Note that in the small-scale realistic Standford2D3D dataset, PGFuse outperforms other methods in all metrics. In [7], the authors argue that due to the limited data in the Stanford2D3D dataset, the transformer-based approach can hardly provide accurate global perception, while the core components of our method are CNN-based and cannot be influenced by this negative effect. As for the Structred3D dataset, we can observe that our approach performs poorly on the Sq Rel metric than the Unifuse and Elite360D, we suspect it might be related to the pre-processing operation. For a fair comparison, we follow the Elite360D to process the data (normalize the data and scale it to 10m), which confuses the model in estimating an accurate depth range.
[…]
@@
Complexity. We show the complexity comparison results in Table II. We can observe that those transformer-based architectures may have fewer parameters and FLOPs. However, the complexity of these approaches is quadric to the input pixels, which limits its application to high-resolution images. In addition, their graphics memory cost is huge, which makes it difficult to use in practice. The other SoTA approach, Bifuse, has the largest parameters and huge FLOPs. And the training time is unbearable for large-scale datasets. Although our scheme has slightly more parameters and FLOPs than the Elite360D, the proposed approach performs better than Elite360D, especially on the RMSE metric, by a significant margin. In addition, we argue that the core fusion component of the Elite360D is attention-based, similar to the transformer-based frameworks, resulting in fewer parameters and FLOPs.

### v5

@@ II-A Monocular 360 Depth Estimation
Panoramic distortion leads to notable performance degradation in perspective image-based depth estimation approaches. To this end, Zioulis etal.et~al. [43] highlighted the need to learn within the 360 domain, introducing the 3D60 dataset. Moreover, to mitigate distortion, researchers have adjusted the sampling positions of convolutional kernels for distortion awareness [15, 36, 8]. Some indoor monocular panoramic depth estimation methods [11, 41, 21] have developed various distortion-aware techniques to manage panoramic distortions, but most of these techniques are computationally inefficient. Recently, advanced Transformer-based architectures [24] have broadened the receptive field, enabling models to extract features from 360 images. However, the traditional self-attention mechanism struggles to handle the complex distortion distribution in 360 images [28]. Enhanced distortion-aware transformer architectures [28, 39, 23, 22, 42, 30], while effective, often come with high computational and memory costs. To simplify models, some approaches [32, 31, 26, 29, 10] advocate for estimating indoor depth from vertically compressed one-dimensional sequences, which leverages gravity alignment properties. However, this compression often fails to capture essential structural geometric information. To achieve a balanced trade-off between performance and efficiency, the mainstream approaches[37, 2, 3] still prefer to develop a more efficient fusion-based framework. However, these fusion methods neglect the reintroduced distortion during the fusion process. Besides, although many of these approaches claim state-of-the-art performance, comparisons are often complicated by variations in datasets, preprocessing methods, and evaluation strategies, making fair comparisons challenging. To this end, Elite360D [3] provides a solid baseline. Therefore, in this paper, we follow the Elite360D method [3] to conduct the experiments. Different from prior fusion frameworks such as UniFuse [17], which rely on ERP–cubemap pairs and simple concatenation strategies, our method adopts dual cubemaps with a single shared-weight encoder and introduces a distortion-aware PanoGabor fusion layer to explicitly correct reintroduced ERP distortions. To our knowledge, this is the first framework that explicitly addresses the problem of reintroduced distortions in fusion, thereby defining a novel paradigm for panoramic depth estimation beyond existing dual-projection fusion-based architectures.
@@ II-B Gabor Filters in Image Processing
The Gabor filter is widely used in texture analysis, object detection, and face recognition due to its robustness to varying angles and scales. Li etal.et~al. [20] suggested enhancing face recognition by learning the Gabor wavelet’s covariance matrix. In fabric defect detection, Chen etal.et~al. [9] introduced the Genetic Algorithm Gabor Faster R-CNN (Faster GG R-CNN), embedding Gabor kernels to reduce texture interference. Andrey etal.et~al. [4] proposed GaborNet, which uses Gabor filters to improve convergence and reduce the complexity of training in image recognition. To extend the receptive fields and enhance the geometric depth cues, we introduce Gabor filters to analyze texture in the frequency domain. Considering the special characteristics of 360 images, we propose a distortion-aware Gabor-based approach to address the challenging distortion issues.
@@ III Methodology
In this section, we introduce PGFuse, whose overall pipeline is illustrated in Fig. 2. The proposed framework is designed as a two-stage distortion-optimization process, in which cubemap-based feature extraction and PanoGabor-based fusion are deployed in complementary domains to address distinct distortion sources. Specifically, equirectangular projection (ERP) suffers from severe non-uniform geometric distortion, particularly near the poles, which greatly hinders the effectiveness of convolutional feature extraction. While PanoGabor is effective in refining distortion-aware frequency responses, it cannot fundamentally remove the intrinsic geometric deformation of ERP. To mitigate this issue, we first project the sphere into the cubemap domain, which decomposes the panorama into locally tangential planes with significantly reduced geometric distortion. This design also allows us to take advantage of perspective-pretrained backbones, thereby improving both training efficiency and feature quality. To further alleviate pixel-missing and discontinuous boundary problems, we adopt a dual-cube strategy, in which two cubemaps—one in the original orientation and one rotated by 45—are jointly exploited [29]. These features are spatially aligned, unified into the ERP format, and subsequently processed by our distortion-aware fusion mechanism.
At this stage, a second type of distortion inevitably arises: projection-induced resampling artifacts and cube-face seams during the unification back into ERP space. To explicitly address this problem, we design PanoGabor filters and embed them into the CS-UFM fusion module. The PanoGabor filters incorporate a latitude-aware distortion representation that approximates the cosine weighting of latitude and modulates frequency-domain responses accordingly. In this way, they enhance depth cues, smooth boundary discontinuities, and ensure global consistency of the ERP representation. Finally, the refined features are fed into a multi-scale distortion-aware module (MS-DAM) [30] to integrate shallow geometric and deep semantic cues, followed by a hierarchical up-sampling decoder that produces the final depth map. Additional spherical gradient constraints are imposed by projecting the ERP depth map onto tangent patches, where Sobel operators are applied to stabilize orientation sensitivity. The following subsections present the details of each component.
@@ III-A Preliminary
The distortion factor $K(\phi)$ ranges from $1$ to $\infty$. At the equator, the projection corresponds to an identity mapping, implying that the distortion is zero at this latitude. To more accurately describe the variation in distortion across latitudes, we introduce a bias term, leading to the following expression for the distortion factor:
@@
Gabor Transform. When applied to image processing, the Gabor transform can provide useful spatial frequency and orientation analysis of the images. By responding to specific spatial frequencies and orientations, the Gabor filters can effectively identify textures and detect edges, which are critical for depth estimation. The standard Gabor transform formula can be denoted as:
@@
[…]
11: Calculate Gabor filter $g\leftarrow\exp\left(-0.5\cdot\left(\frac{x_{rot}^{2}+y_{rot}^{2}}{\sigma^{2}}\right)\right)\cdot\cos(f\cdot x_{rot}+\psi);gg$$\leftarrow\frac{g}{2\cdot\pi\cdot\sigma^{2}}$;
@@ III-B Distortion-aware PanoGabor Filters (PGF)
where the range of $d_{c}$ is $0,+\infty)$. The scale factor mirrors the increasing distortion in equirectangular projections towards the poles. Subsequently, as detailed in Algorithm 1, we employ the defined distortion coefficient to control the generation of distortion-aware PanoGabor filters, thereby creating a bank of PanoGabor filters.
@@
Figure 5: Illustration of the spherical tangent sampling. A representation that the pixels are projected from the pixel coordinates to the spherical coordinates to sample tangent points. b indicates that the calculated sampling points are projected back to the pixel coordinates.
@@ III-C Spherical Gradient Constraint
Create sampling mesh grid. Given a depth map $D\in\mathbb{R}^{H\times W}$, we first define the sampling grid. For each pixel in the depth map, we sample a 3×3 tangent patch. Let X=[1,0,1]X=[-1,0,1] and Y=[1,0,1]Y=[-1,0,1] represent the sampling kernels for the $x$- and $y$-directions, respectively. We define the 2D grid mesh $M$ (Fig. 5left) as:
@@ III-D Channel-wise and Spatial-wise Unidirectional Fusion Module (CS-UFM)
The fusion module is designed to integrate two sets of ERP features, addressing reintroduced distortion, inconsistencies in cubemap boundaries, and pixel loss issues. This module (as illustrated in Fig. 6) first concatenates the two groups of input ERP features and uses a residual block to achieve an initial refinement. In particular, the block comprises a 1x1 convolution that compresses the channels to match the features’ height, facilitating the application of the PanoGabor layer. Following the 1×1 convolution layer, a 3×3 PanoGabor layer is used to process the features, aiming for distortion awareness and depth cues enhancement.
@@ IV Experiment
We conduct all experiments on a single NVIDIA GeForce RTX 3090 GPU, implementing our approach using the PyTorch framework. We use the Adam optimizer [19] with an initialized learning rate of $1\times 10^{-4}$ to train our model. We keep the default settings for other optimizer hyperparameters. During the training period, we use a batch size of 8 and employ left-right flipping, panoramic horizontal rotation, and luminance change for data augmentation, as previously proposed in works [18, 28]. The backbones we employed are all pre-trained on ImageNet-1K.
Dataset. Two real-world datasets for experimental validation: Stanford 2D-3D [5], and Matterport3D [6]. The depth maps in Stanford 2D-3D are constructed from reconstructed 3D models, and some depths are inaccurate. Besides, the bottom and top regions of the RGB images are missing. We follow the official splits and get 1,040 RGB-D pairs for training and 343 for testing. The Matterport3D dataset provides raw depth whose top and bottom depths are missing. For the RGB panoramas in Matterport3D, we follow the Unifuse method [17] to fill the missing regions. Finally, we get 7829 pairs for training and 2014 pairs for testing.
One virtue dataset: Structured3D [40]. The virtue dataset was first proposed to explore the layout estimation task. Recently, some researchers have preferred to evaluate the depth estimation models on the dataset for its high-quality annotations. We follow the official splits that we employ the frame scene_0000 to frame scene_02999 for training and scene_03250 to scene_03499 for testing. We especially follow the official guidelines and remove invalid cases.
@@
Figure 7: Qualitative comparison results. We exhibit the predicted depth maps of the Unifuse [18], Elite360D [3], and ours. All depth maps adopt a unified visualization method.
Figure 8: Visualization of the point cloud generated with the ground truth (GT) and the predicted depth map (Ours, Elite360D and UniFuse). The areas marked in red rectangles indicate that our approach gains superior performance (dense and high-quality point cloud).
@@ IV-B Comparison Analysis
Quantitative Results. We exhibit the quantitative comparison results in Table LABEL:tab:quanti, which demonstrate that our model outperforms all other state-of-the-art solutions on most metrics. Our PGFuse method significantly outperforms the compared methods. Specifically, on the real-world Matterport3D and Stanford2D3D datasets, our approach demonstrates notable improvements in the important RMSE and $\delta_{1}$ metrics. Additionally, it achieves comparable results on the virtual Structured3D dataset, showcasing its robustness and effectiveness across various environments. These results highlight the efficacy of PGFuse in monocular 360 depth estimation across diverse and challenging datasets. For the Matterport3D dataset, PGFuse consistently outperforms other methods on the accuracy metrics with stricter thresholds $\delta_{1}$ and $\delta_{2}$, as well as on RMSE and Abs Rel. However, on the metric $\delta_{3}$, PGFuse exhibits slightly lower performance than that of Elite360D [3] and UniFuse [17]. Since the $\delta_{3}$ metric uses the loosest threshold and is almost saturated for near- and mid-range depths for all methods in comparison, it is relatively insensitive to the improvements brought by our design, which mainly targets local details. Note that in the small-scale, realistic Stanford2D3D dataset, PGFuse outperforms other methods in all metrics. In [3], the authors argue that due to the limited data in the Stanford2D3D dataset, the transformer-based approach can hardly provide accurate global perception, while the core components of our method are CNN-based and cannot be influenced by this negative effect. As for the Structured3D dataset, our method achieves superior performance on most metrics, but the Sq Rel value is noticeably higher than that of UniFuse and Elite360D. We attribute this primarily to the depth preprocessing protocol. Specifically, for a fair comparison, we follow Elite360D by normalizing the depth maps and rescaling them to a maximum range of 10 m. This global rescaling changes the original depth distribution and introduces a slight shift in depth scale. Since PanoGabor estimates depth from Gabor-based learning of local texture patterns, a scale mismatch between the normalized depths and the original scene geometry can disturb this texture–depth mapping. Under this preprocessing protocol, such shifts in the depth scale tend to produce a modest increase in the Sq Rel metric.
Qualitative Results. We exhibit the qualitative comparison results with the current SoTA UniFuse [18] and Elite360D [3] method. As shown in Fig. 7, we can observe that UniFuse can effectively remove the negative effect of the panoramic distortions, but it still suffers from losing detailed information. This degradation is due to the projection format of the cubemap that loses 25% of pixels when converted to ERPs. In contrast, Elite360D alleviates such information loss and provides relatively sharper depth boundaries, yet it fails to recover fine-grained structural cues in complex scenarios. Our method preserves a competitive depth range and maintains accurate details. This preservation of details contributes to superior performance in RMSE metrics. In addition, from Fig. 8 our method shows remarkable results even in the regions where the ground truth is wrong.
@@
Complexity. We show the complexity comparison results in Table II. We can observe that those transformer-based architectures may have fewer parameters and FLOPs. However, the complexity of these approaches is quadratic to the input pixels, which limits their application to high-resolution images. In addition, their graphics memory cost is huge, which makes them difficult to use in practice. The other SoTA approach, Bifuse, has the largest parameters and huge FLOPs, and the training time is unbearable for large-scale datasets. Although our scheme has slightly more parameters and FLOPs than Elite360D, the proposed approach performs better than Elite360D, especially on the RMSE metric, by a significant margin. In addition, we argue that the core fusion component of Elite360D is attention-based, similar to transformer-based frameworks, resulting in fewer parameters and FLOPs. To further evaluate computational efficiency, we benchmark both inference speed and memory on a single NVIDIA RTX 3090 (24 GB) under the same setting (batch size = 8). Our method achieves 40 FPS with 5770 MB GPU memory, whereas Elite360D attains 21 FPS with 7728 MB. This corresponds to 25 ms/image for our method versus 48 ms/image for Elite360D, demonstrating balanced efficiency by delivering both higher throughput and lower memory consumption under identical hardware conditions.
@@
IV-C Ablation Study
To demonstrate the effectiveness of the main proposed components, we follow Unifuse [18] and Elite360D [3] to design the ablation study on the Matterport3D dataset [7]. All the models have the same conditions and are trained to converge. As illustrated in Fig. 2, we remove the fusion stage from the full framework to serve as the baseline for conducting the ablation study. The experimental results of the baseline are presented as ‘a’ in Table LABEL:tab:ablation. We further add the proposed CS-UFM module, a version without the PanoGabor layer embedded as ‘b’. Index ‘c’ refers to the use of the full CS-UFM module. Index ‘d’ indicates that we employ the proposed spherical gradient constraint to stabilize the orientation sensitivity of the Gabor filters. Additionally, in ‘e’, all PanoGabor layers are replaced with normal Gabor filters to further verify the effectiveness of the designed PanoGabor.